OpenStudy (anonymous):

Solve the following system of equations using the substitution method. F = C F = 36 + (9/5)C

OpenStudy (anonymous):

F=36+(9/5)F F-9F/5=36 (5F-9F)/5=36 -4F=36*5 F=180/-4 F=-45 F=-4,C=-45